Conservation management for hairy-nosed wombats also includes a captive breeding program. In the captive population of 77 wombats, you determine that the females in the population produce 20 surviving joeys each year. On average, 15 wombats in the captive population die every year. Based on these numbers, and given unlimited space and resources, how many wombats would be in your captive population in five years? Ten years? Twenty years? Show all your work.
Answer #1

Population of wombats in year 1 = 77+20-15 = 82.

Population of wombats in year 2 = 82+20-15 = 87

Therefore there is a net increase of 5 wombats each year

Therefore ,number of new wombats after 5 years = 5x5 =25

Therefore ,number of new wombats after 10 years = 5x10 =50

Therefore ,number of new wombats after 10 years = 5x20 =100.

N0: wombats after 5 years = 77+ 25 = 102

No: wombats after 10 years = 77 +50 = 127

No: wombats after 20 year = 77+100 =177